OpenStudy (anonymous):

Use elimination to solve. X-3y=10 X+2y=15

OpenStudy (luigi0210):

multiply the top by a negative

OpenStudy (anonymous):

What negative?

OpenStudy (kenljw):

Subtract second eq from first -5y = -5 : y= 1 : substitute in either eq x+2(1) = 15 : x =13 x-3(1) =13 : checks

OpenStudy (kenljw):

x-3(1) =10 : checks

OpenStudy (luigi0210):

We're using elimination not substitution.

OpenStudy (anonymous):

Yeah eliminatin

OpenStudy (kenljw):

Use elimination first by subtract eq 2 from eq 1, then subtitution

OpenStudy (luigi0210):